Assuming a discount rate of 6.6 percent, what is the present value of the following stream of uneven cash flows over the next 10 years?

Year 1 $26,728
Year 2 $19,363
Year 3 $11,392
Year 4 $17,496
Year 5 $11,100
Year 6 $12,041
Year 7 $10,304
Year 8 $27,646
Year 9 $26,151
Year 10 $16,536
Answer #1

012 - fx =SUM(D2:011) AB 1 year Cash flows pv@6.6% Present value 11 $ 26,728.00 0.9381 $ 25,073.17 21 $ 19,363.00 0.8800 $ 17

- fx =SUM(D2:011) B 012 A 1 year 2 1 D Present value =B2*C2 32 4 3 65 76 87 98 10 9 11 10 12 Cash flows 26728 19363 11392 174